OpenStudy (omarbirjas):

Will Medal 3surd((x+2)^3,5)+3=27 copy and paste into wolframalpha.com and you'll see the equation.

OpenStudy (omarbirjas):

\[3\sqrt[5]{(x+2)^3}+3=27\]

OpenStudy (acxbox22):

are you solving for x?

OpenStudy (omarbirjas):

Yes

OpenStudy (acxbox22):

firstly divide both sides by 3

OpenStudy (omarbirjas):

we get this right? \[(\sqrt[5]{(x+2)^3})/3+1=9\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i think

OpenStudy (acxbox22):